OpenStudy (anonymous):

An IT user support service has monitored the daily number of phone calls requesting for IT services, No of calls 0 1 2 3 4 5 6 Probability K 0.15 0.3 0.2 K K 0.05 i) determine the value of K. ii) over 3 days, what is the probability of IT services being requested more than twice on each day?

OpenStudy (anonymous):

k = .1

OpenStudy (anonymous):

how you get that?

OpenStudy (anonymous):

P = 1

OpenStudy (anonymous):

the sum of that must be 1.

OpenStudy (anonymous):

(0xK)+(1x0.15)+(2x0.3) + (3x0.2)+(4xK)+(5xK)+(6x0.05) = 1 K = 0.18

OpenStudy (anonymous):

is this correct?

OpenStudy (anonymous):

ii) (1-p(0)-p(1))³

OpenStudy (anonymous):

no just the sum of prob must be one(K + 0.15 +0.3 +0.2 +K + K +0.05)

OpenStudy (anonymous):

so its K + 0.15 +0.3 +0.2 +K + K +0.05 = 1 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so k = 0.1 . ii) (1-p(0)-p(1))³

OpenStudy (anonymous):

@RaphaelFilgueiras That's clever =)

OpenStudy (anonymous):

i dont know how to get p(0) and p(1)

OpenStudy (anonymous):

= [1 - ( .1 + .15) ] = .44

OpenStudy (anonymous):

thank you so much!